[figure: 9.7 cm]

FAST BOLOMETRIC RESPONSE OF BULK La _{0.7}Sr_{0.3}MnO_{3} ELECTROCERAMICS

Yu.M. Nikolaenko^{1}, I.S. Maksimov^{1}, Yu.V. Medvedev^{1}, A.N. Ulyanov^{1}, A.M. Grishin^{2}

1Donetsk Physical-Technical Institute of the National Academy of Sciences of Ukraine,

72 R.Luxemburg, Donetsk, 340114, Ukraine 2Departament of Condensed Matter Physics, Royal Institute of Technology, S-100 44, Stockholm, Sweden

We report on the performance of a microwave electroceramic bolometer (MECB) of hybrid La_{0.7}Sr_{0.3}MnO_{3}/Al_{2}O_{3} (0.224 mm^{3}) structure. The estimated thermal resistance of the ceramic manganate - single crystal sapphire interface is about 500 K/W at room temperature. This resistance is the main thermal barrier in the heat sink system and has been found to be slightly dependent on the temperature. When compared with high-T_{C} superconducting bolometers, La_{0.7}Sr_{0.3}MnO_{3} MECB works in a wide temperature range 77 K to 330 K excluding the narrow temperature interval at the metal-dielectric phase transition (T = 230 K). The MECB sensitivity and time constant at room temperature have been found to be 0.1 V/W and 100 ms correspondingly. To improve bolometer performance the point contact La_{0.7}Sr_{0.3}MnO_{3} MECB has been fabricated by break-junction technique. The optimization of MECB design brought to considerable improvement of basic bolometer characteristics. Point contact MECB exhibits linear I-V characteristic, microwave sensitivity was about 0.3 V/W and time constant was less than 100 ns.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Igor Maksimov

Address for correspondence : Donetsk Physical and Technical Institute of NASU 72, str. R.Luxemburg 340114 Donetsk Ukraine FAX: (0622) 55-74-62

Email address : maksimov@kinetic.ac.donetsk.ua